OpenStudy (anonymous):

5. Fill in the blank with the correct number to convert 0 = x2 - 4x + 8 into transformational form 0 = (x - ___)2 + 4. A. 4 B. 8 C. 16 D. 2

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yeah how did you get B i have to show work @FATBOI2K12

OpenStudy (anonymous):

\[0=x^2-4x+8=x^2-4x+4+4=(x-2)^2+4\]

OpenStudy (anonymous):

can you help me with a few more @surjithayer

OpenStudy (anonymous):